# DF40: Toward Next-Generation Deepfake Detection

### v1

@@ 6 Open Questions and Potential Topics for Future Research
(3) Previous research [93; 92] classified deepfake types at the instance level. For example, if four deepfakes are applied for training, they will be regarded as four distinct deepfakes. However, there are so many forgery techniques available that it is hard to elaborate on them all. Question-3: can we classify deepfakes based on FS, FR, EFS, and FE? It is still questionable since different forgeries can share similar forgery patterns, $e.g.$, generative artifacts (see Fig. 8). So, it would be possible to define a "metric" to classify different forgeries into different classes, aiming to learn effectively specific forgery features and common features jointly (following the framework of UCF [93]).
(4) In our research, it can be seen that CLIP-large can achieve obvious advantages over other models (including CLIP-base). We perform t-SNE (Fig. 7) and logit distribution (Fig. 6) for deeper analysis. We found that CLIP-large can learn a more robust distribution for the real face than CLIP-base. So, Question-4: can we further "inflate" this CLIP-large model into a video model for detecting deepfake videos? Or similarly, Question-5: can we leverage the power of the current SoTA multi-modality technique to develop a better real face representation for deepfake detection?
7 Conclusions, Board Impacts, Limitations, and Dataset Usage
(1) Conclusions: We have developed DF40, a highly diverse and groundbreaking benchmark, comprising 40 distinct deepfake techniques to support the detection. Leveraging DF40, we conducted over 2,000+ evaluations using 7 representative detectors under 4 standard evaluation protocols, creating 12 new findings and 5 open questions for future works. We hope the proposed DF40 could revolutionize the whole field for the next generation.
(2) Board Impacts: DF40 offers high-quality and realistic deepfake techniques, facilitating the detection of today’s real-world deepfakes. Also, our benchmark assists in safeguarding societal trust and promoting the responsible use of such technology.
@@ 7 Conclusions, Board Impacts, Limitations, and Dataset Usage
(3) Limitations: One limitation is the lack of comprehensive analysis for video-level detectors. Actually, we provide evaluations using video models, $e.g.$, I3D [9] in the supplementary. However, we have not delved deeply into discussing and analyzing certain issues, $e.g.$, assessing and visualizing whether video models can effectively capture both temporal and spatial artifacts. In future work, we plan to broaden our benchmark’s scope and address these concerns in detail.
(4) Dataset Usage: The proposed DF40 dataset can be incorporated seamlessly into the codebase of DeepfakeBench [94], where the user can use the existing more than 30 detection methods to train or test on our DF40, making it convenient to use. We will release the dataset on GitHub.
8 Content Structure in Supplementary
We have organized additional important content in the supplementary due to the limited space. We present a brief outline of the content structure of the supplementary to facilitate readers to find the corresponding content, as follows:
  • Section A.1: Dataset Generation Methods & Original Data:
    • Section A.1.1: Brief Introduction of Generation Methods;
    • Section A.1.2: Details of 40 Implemented Synthesis Methods;
    • Section A.1.3: Formulation of Manipulation Methods;
    • Section A.1.4: Details of Original Datasets;
  • Section A.2: Introduction of the Used Detection Methods:
  • Section A.3: Experimental Setup & Full Results:
    • Section A.3.1: Experimental Setup and Details;
    • Section A.3.2: Full Experimental Results and Further Discussion;
  • Section A.4 Additional information on Dataset Publication:
    • Section A.4.1: Descriptions of Links;
    • Section A.4.2: Descriptions of the Hosting Platform and Access to the Dataset;
    • Section A.4.3: Descriptions of licensing.
9 Takeaways
In this section, we very briefly outline several key takeaways to encapsulate our contributions and conclusions:
  • (1) The dataset is pivotal in addressing the generalization problem in deepfake detection. To this end, we build a highly diverse deepfake dataset incorporating 40 distinct deepfake techniques, including the most recent ones. The visual example can be seen in Fig. 12 (FS), Fig. 13, Fig. 14 (EFS), and Fig. 15 (FE and other novel methods);
  • (2) Data domain and forgery method collectively determine the final detection results (see the causal graph Fig. 5 for intuitive understanding);
  • (3) Many recent face-swapping methods ($e.g.$, Inswapper [33], SimSwap [13]) do not involve a blending process but generate all content directly (including the background). See the artifacts between SimSwap (FS) and SD1.5 (EFS) in Fig. 8.
  • (4) Blending is not all you need to detect deepfakes, even face-swapping forgeries (Sec. 5.1);
  • (5) CLIP-large is the most powerful baseline model due to its notable ability to learn robust real-face distribution (see Sec. 5.3);
  • (6) The model trained on the face domain (our dataset) can also be somehow applied to the non-face domain (pure AIGC) fakes (see Sec. 5.5);
  • (7) There is an urgent need to develop an effective incremental learning framework aiming to produce improved results, creating "1+1>2" results when combining many different forgeries together for training (see Question-2).
@@ 9 Takeaways
We would like to express our profound gratitude to all the individuals (all collaborators) who have contributed to the successful completion of this work. Our heartfelt appreciation goes out to Yue Han and Chengming Xu from Tencent Youtu Lab, as well as Wentang Song from Shenzhen University, for their invaluable support in code implementation and reproduction.
Furthermore, we extend our gratitude to Ke Sun from Xiamen University, Mingli Zhu from the Chinese University of Hong Kong, Shenzhen (CUHK-Shenzhen), Shaokui Wei from CUHK-Shenzhen, Mingda Zhang from CUHK-Shenzhen, and Yuhao Luo from CUHK-Shenzhen for their insightful discussions and contributions to this work.
Our sincere thanks go to the code developers of DeepfakeBench for their easy-to-use codebase and unified training/testing framework. This includes Xinghang Yuan (previously in CUHK-Shenzhen), Yize Chen (CUHK-Shenzhen), Kangran Zhao (CUHK-Shenzhen), and Jikang Cheng (Wuhan University). Once again, we thank everyone involved for their dedication and efforts in making this big project successful.
@@
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] Our contributions are in Lines 53-59 of the "introduction" section. We claim the scope of deepfake detection in Lines 36-39 of the "introduction" section, and we further clarify our research scope in Lines 105-108 of the "DF40 benchmark" section.
@@
      Did you discuss any potential negative societal impacts of your work? [No] Our work is to detect those potentially "dangerous" deepfake techniques.
@@
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] We have written detailed instructions to reproduce the main experimental results in the supplementary.
@@
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] We have specified all these details in the supplementary.
@@
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] We have shown these details in the supplementary.
@@
      Did you mention the license of the assets? [Yes] We provide the license in the supplementary.
@@
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] All real people in our work are directly cited from the previous/existing works.
@@
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] All real people in our work are directly cited from the previous/existing works.
@@ Appendix A Appendix
A.1 Dataset Generation Methods & Original Data
A.1.1 Brief Introduction of Generation Methods
@@
(1) Face-swapping: This paper classifies the face-swapping technique into two domains: DF-family and FS-family. (i) DF-family involves creating a mask around the facial region (some even include the neck [50]) and blending the generated deepfake face back into the background image using that mask. Most existing and famous face forgery datasets, such as FF-DF [67], Celeb-DF [47] and DFDC [20], belong to this line. (ii) FS-family methods represent another significant category in face-swapping deepfake generation. These methods typically involve the use of an identity-background encoder. It disentangles a face image’s identity and background information during encoding. Notably, these methods directly generate all content, even the background. Much recent face-swapping research works [90; 13] and popular software ($e.g.$, Roop) are within this line.
(2) Face-reenactment: Generally, this technique can be used to modify source faces, imitating the actions or expressions of another face. Differing from face-swapping, face-reenactment techniques are rarely considered in existing datasets. Two commonly used reenactment-based forgeries are Face2Face [77] and NeuralTextures [76]. These two forgeries are implemented in the FF++ dataset. Face2Face employs pairs of original and target faces, using key facial points to generate varied expressions, while NeuralTexture uses rendered images from a 3D face model to migrate expressions. Although these forgeries achieve more realistic visual synthetic results compared to their face-swapping counterparts in FF++, Due to the amazingly rapid development of the AIGC technologies ($e.g.$, Digital Human), these relatively old-fashioned methods cannot represent the modern’ SoTA reenactment methods. Our DF40 implements 13 face-reenactment methods in total, including the classical animation [72; 73], SoTA’s audio-based driven methods [97; 60], image-based driven methods [86; 31; 7]. We also include the well-known best face generation technique, HeyGen, in our dataset for evaluation.
@@
A.1.2 Details of Implemented Forgery Methods in DF40
@@ 40. StyleCLIP [58]
A.1.3 Formulation of Manipulation Methods
To guarantee the diversity of deepfake approaches in the proposed DF40, we introduce and implement 40 distinct deepfake techniques, which are listed in Tab. 2. A detailed description of each deepfake method is provided in the supplementary. They are selected according to perspectives of modeling types, conditional sources, forgery effects, and functions. Formally, we denote $x_{t}(i_{t},a_{t},b_{t})$ as the target subject to be manipulated, which possesses attributes ($i$: person identity, $a$: identity-agnostic content, $b$: external attributes) that can uniquely determine itself, while the source $x_{s}(i_{s},a_{s},b_{s})$ is regarded as the conditional media ($c$), driving the target to change either identity or attributes or even both. We use $f_{sw},f_{re},f_{sy},f_{ed}$ to represent four types of deepfakes in our work: FS, FR, EFS, and FE, respectively. Fig. 1 provides some visual examples for each category in DF40, and Fig. 16 generally illustrates them from the method perspective. (i) FS Approach: From Fig. 16(a), FS aims to replace the content of $x_{t}$ with that of $x_{s}$ preserving the identity $i_{s}$. Formally, $\tilde{x}_{t}(\tilde{i}_{s},a_{t},b_{t})$ only swaps identity $i$ from the source $x_{s}$ to the target $x_{t}$, and the identity-agnostic content $a$ are preserved. The swapping process can be expressed as:
@@ A.1.3 Formulation of Manipulation Methods
A.1.4 Details of Original Datasets
@@
The Celeb-DF dataset is a large-scale dataset containing both real and manipulated celebrity videos, aimed at facilitating the development of deepfake detection algorithms. CDF dataset has two different versions: CDF-v1 and CDF-v2. Mostly, CDF-v2 is the default version of CDF. Here, we also adopt the CDF-v2 version. It consists of more than 5900 videos, with 563 real videos and over 5300 deepfake videos generated using a modified version of the DeepFakes approach. The dataset is designed to evaluate the performance of deepfake detection methods in real-world scenarios. The Celeb-DF dataset is led by the State University of New York at Buffalo. The dataset can be accessed through the following link: https://github.com/yuezunli/celeb-deepfakeforensics. The real human data sources for this dataset are not explicitly mentioned. The copyright information can be found at https://docs.google.com/forms/d/e/1FAIpQLScoXint8ndZXyJi2Rcy4MvDHkkZLyBFKN43lTeyiG88wrG0rA/viewform?pli=1.
@@
A.2 Introduction of the Used Detection Models
@@
A.3 Experimental Setup and Full Experimental Results
A.3.1 Experimental Setup
@@
A.3.2 Full Experimental Results and Discussion

### v2

@@ 4.4 Open Questions and Potential Topics for Future Research
(3) Regrading Forgery Type: Previous research [89; 88] classified deepfake types at the instance level. For example, if four deepfakes are applied for training, they will be regarded as four distinct deepfakes. However, there are so many forgery techniques available that it is hard to elaborate on them all. So, Question-3: can we classify deepfakes based on FS, FR, EFS, and FE?
(4) Domain-Invariant Detector: As shown in our previous discussion, we have found that many factors (such as resolution and fake region) can impede the model from learning domain-invariant features. Thus, Question-4: How can we develop an invariant detector?
5 Conclusions, Board Impacts, and Limitations
@@
(1) Conclusions: We have developed DF40, a highly diverse and groundbreaking benchmark, comprising 40 distinct deepfake techniques to support the detection. Leveraging DF40, we conducted over 2,000+ evaluations using 8 representative detectors under 4 standard evaluation protocols, creating 7 new findings and 4 open questions for future works. We hope the proposed DF40 could revolutionize the whole field for the next generation. (2) Board Impacts: DF40 offers high-quality and realistic deepfake techniques, facilitating the detection of today’s real-world deepfakes. Also, our benchmark assists in safeguarding societal trust and promoting the responsible use of such technology. (3) Limitations: One limitation is the lack of comprehensive analysis for video-level detectors. Actually, we provide evaluations using video models, $e.g.$, I3D [7] in the Appendix. However, we have not delved deeply into discussing and analyzing certain issues, $e.g.$, assessing and visualizing whether video models can effectively capture both temporal and spatial artifacts. In future work, we plan to broaden our benchmark’s scope and address these concerns in detail.
@@
This work was supported in part by the Natural Science Foundation of China (No. 62202014, 62332002, 62425101), Shenzhen Basic Research Program (No.JCYJ20220813151736001).
We would like to express our profound gratitude to all the individuals (all collaborators) who have contributed to the successful completion of this work. Our heartfelt appreciation goes out to Yue Han and Chengming Xu from Tencent Youtu Lab, as well as Wentang Song from Shenzhen University, for their invaluable support in code implementation and reproduction. Furthermore, we extend our gratitude to Ke Sun from Xiamen University, Mingli Zhu from the Chinese University of Hong Kong, Shenzhen (CUHK-Shenzhen), Shaokui Wei from CUHK-Shenzhen, Mingda Zhang from CUHK-Shenzhen, and Yuhao Luo from CUHK-Shenzhen for their insightful discussions and contributions to this work. Our sincere thanks go to the code developers of DeepfakeBench for their easy-to-use codebase and unified training/testing framework. This includes Prof. Baoyuan Wu (project leader), Xinghang Yuan (previously in CUHK-Shenzhen), Yize Chen (CUHK-Shenzhen), Kangran Zhao (CUHK-Shenzhen), and Jikang Cheng (Wuhan University). Once again, we thank everyone involved for their dedication and efforts in making this big project successful.
@@
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] Our contributions are briefly summarized within the "introduction" section, and we further clarify our research scope within the "DF40 benchmark" section.
@@
      Did you discuss any potential negative societal impacts of your work? [Yes] Yes, we discuss the potential negative impacts and also propose an approach for controlled access (see Sec. 5 and Sec. LABEL:2.2 in the Appendix.
@@
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] We have written detailed instructions to reproduce the main experimental results in the Appendix.
@@
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] We have specified all these details in the Appendix.
@@
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] We have shown these details in the Appendix.
@@
      Did you mention the license of the assets? [Yes] We provide the license in the Appendix.
@@
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] All real people in our work are directly cited from the previous/existing research works.
@@
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] All real people in our work are directly cited from the previous/existing research works.
@@ Appendix A Appendix
A.1 Content Structure in Appendix
We have organized additional important content in the Appendix due to the limited space. We present a brief outline of the content structure of the Appendix to facilitate readers to find the corresponding content, as follows:
  • Section A.2: Dataset Generation Methods & Original Data:
    • Section A.2.1: Brief Introduction of Generation Methods;
    • Section A.2.2: Details of 40 Implemented Synthesis Methods;
    • Section A.2.3: Formulation of Manipulation Methods;
    • Section A.2.4: Rationale of Fine-tuning EFS Methods;
    • Section A.2.5: Details of Original Datasets;
  • Section A.3: Introduction of the Used Detection Methods:
  • Section A.4: Experimental Setup & Full Results:
    • Section A.4.1: Experimental Setup and Details;
    • Section A.4.2: Full Experimental Results and Further Discussion;
  • Section A.5: Further Analysis Results:
    • Section A.5.1: More Results of Different Fake Regions;
    • Section A.5.2: Generalization to Non-face Domain AIGCs;
    • Section A.5.3: Generalization to External/Previous Deepfake Datasets;
    • Section A.5.4: Comparative Analysis of CLIP-base and CLIP-large for Deepfake Detection;
    • Section A.5.5: Analysis of "Anomalous Values" in Experiments;
    • Section A.5.6: Comparison of Video Model and Image Model;
    • Section A.5.7: Visualizations of the Real People Features;
    • Section A.5.8: Artifacts of deepfake forgeries in Frequency;
  • Section A.6 Additional information on Dataset Publication:
    • Section A.6.1: Hosting Platform and Links;
    • Section A.6.2: Controlled Access;
    • Section A.6.3: Discrimination, Bias, and Fairness.
A.2 Dataset Generation Methods & Original Data
A.2.1 Brief Introduction of Generation Methods
@@
(1) Face-swapping: This paper classifies the face-swapping technique into two domains: DF-family and FS-family. (i) DF-family involves creating a mask around the facial region (some even include the neck [46]) and blending the generated deepfake face back into the background image using that mask. Most existing and famous face forgery datasets, such as FF-DF [62], Celeb-DF [42] and DFDC [16], belong to this line. (ii) FS-family methods represent another significant category in face-swapping deepfake generation. These methods typically involve the use of an identity-background encoder. It disentangles a face image’s identity and background information during encoding. Notably, these methods directly generate all content, even the background. Much recent face-swapping research works [85; 9] and popular software ($e.g.$, Roop [60]) are within this line.
(2) Face-reenactment: Generally, this technique can be used to modify source faces, imitating the actions or expressions of another face. Differing from face-swapping, face-reenactment techniques are rarely considered in existing datasets. Two commonly used reenactment-based forgeries are Face2Face [72] and NeuralTextures [71]. These two forgeries are implemented in the FF++ dataset. Face2Face employs pairs of original and target faces, using key facial points to generate varied expressions, while NeuralTexture uses rendered images from a 3D face model to migrate expressions. Although these forgeries achieve more realistic visual synthetic results compared to their face-swapping counterparts in FF++, Due to the amazingly rapid development of the AIGC technologies ($e.g.$, Digital Human), these relatively old-fashioned methods cannot represent the modern’ SoTA reenactment methods. Our DF40 implements 13 face-reenactment methods in total, including the classical animation [67; 68], SoTA’s audio-based driven methods [92; 55], image-based driven methods [82; 26; 5]. We also include the well-known best face generation technique, HeyGen, in our dataset for evaluation.
@@
A.2.2 Details of Implemented Forgery Methods in DF40
@@ 40. StyleCLIP [53]
A.2.3 Formulation of Manipulation Methods
To guarantee the diversity of deepfake approaches in the proposed DF40, we introduce and implement 40 distinct deepfake techniques, which are listed in Tab. 2. A detailed description of each deepfake method is provided in the Appendix. They are selected according to perspectives of modeling types, conditional sources, forgery effects, and functions. Formally, we denote $x_{t}(i_{t},a_{t},b_{t})$ as the target subject to be manipulated, which possesses attributes ($i$: person identity, $a$: identity-agnostic content, $b$: external attributes) that can uniquely determine itself, while the source $x_{s}(i_{s},a_{s},b_{s})$ is regarded as the conditional media ($c$), driving the target to change either identity or attributes or even both. We use $f_{sw},f_{re},f_{sy},f_{ed}$ to represent four types of deepfakes in our work: FS, FR, EFS, and FE, respectively. Fig. 1 provides some visual examples for each category in DF40, and Fig. 14 generally illustrates them from the method perspective. (i) FS Approach: From Fig. 14(a), FS aims to replace the content of $x_{t}$ with that of $x_{s}$ preserving the identity $i_{s}$. Formally, $\tilde{x}_{t}(\tilde{i}_{s},a_{t},b_{t})$ only swaps identity $i$ from the source $x_{s}$ to the target $x_{t}$, and the identity-agnostic content $a$ are preserved. The swapping process can be expressed as:
@@ A.2.3 Formulation of Manipulation Methods
A.2.4 Rationale of Fine-tuning EFS Methods
Why did we fine-tune the generative model of EFS? The primary motivation is to create an aligned data domain and fake methods for evaluation. We have two main reasons for pursuing this: (1) Alignment with previous evaluation protocols: Most previous deepfake detection studies have followed a widely-used protocol for selecting state-of-the-art detectors: training models on four deepfake methods from FF++ and testing them on one fake method from CDF. This led us to question whether we could expand this approach by implementing more fake methods on FF++ (beyond the original four) and CDF (testing on more than just one specific method). This consideration motivated us to choose FF++ and CDF as the real data sources for generating corresponding fake data. (2) Isolating the individual effects on detection results: Simply following the protocol of training on FF++ and testing on CDF would not allow us to isolate the individual influences of data domains and fake methods on the final detection results. Additionally, we noticed that many previous studies have performed cross-manipulation evaluations—training on one specific method and testing on others within the same data domain (FF++). However, the limited diversity of fake methods in previous datasets restricted cross-manipulation evaluations to just 4-5 different methods, most of which were face-swapping techniques. This limitation motivated us to create a significantly more diverse dataset to enable more comprehensive cross-manipulation evaluations.
How did we fine-tune the generative model of EFS? Please note the fine-tuning details for each EFS method can vary. In Algorithm.3, we provide a general understanding of the fine-tuning process. In details, we also release the specific fine-tuning code for reproduction (https://github.com/YZY-stack/DF40/tree/main/EFS_finetune_code/diffusion_based).
A.2.5 Details of Original Datasets
@@
The Celeb-DF dataset is a large-scale dataset containing both real and manipulated celebrity videos, aimed at facilitating the development of deepfake detection algorithms. CDF dataset has two different versions: CDF-v1 and CDF-v2. Mostly, CDF-v2 is the default version of CDF. Here, we also adopt the CDF-v2 version. It consists of more than 5900 videos, with 563 real videos and over 5300 deepfake videos generated using a modified version of the DeepFakes approach. The dataset is designed to evaluate the performance of deepfake detection methods in real-world scenarios. The Celeb-DF dataset is led by the State University of New York at Buffalo. Specifically, the dataset can be accessed through the following official website: https://github.com/yuezunli/celeb-deepfakeforensics. The real human data sources for this dataset are not explicitly mentioned. The copyright information can be found at https://docs.google.com/forms/d/e/1FAIpQLScoXint8ndZXyJi2Rcy4MvDHkkZLyBFKN43lTeyiG88wrG0rA/viewform?pli=1.
@@
A.3 Introduction of the Used Detection Models
@@
A.4 Experimental Setup and Full Experimental Results
A.4.1 Experimental Setup
Training Config CLIP Other Detectors
Image Size 224, 224 256, 256
Weight Initialization CLIP Pre-trained ImageNet Pre-trained
Optimizer Adam Adam
Base Learning Rate 1e-5 2e-4
Weight Decay 5e-4 5e-4
Optimizer Momentum B1, B2=0.9, 0.999 B1, B2=0.9, 0.999
Batch Size 32 32
Training Epochs 15 15
Learning Rate Schedule None, Constant None, Constant
Flip Probability 0.5 0.5
Rotate Probability 0.5 0.5
Rotate Limit [-10, 10] [-10, 10]
Blur Probability 0.5 0.5
Blur Limit [3, 7] [3, 7]
Brightness Probability 0.5 0.5
Brightness Limit [-0.1, 0.1] [-0.1, 0.1]
Contrast Limit [-0.1, 0.1] [-0.1, 0.1]
Quality Lower 40 40
Quality Upper 100 100
@@ A.4.1 Experimental Setup
A.4.2 Full Experimental Results and Discussion